# Mixed Membership Estimation for Weighted Networks

### v2

@@ 4 Asymptotic Consistency
In this section, we aim at answering Q (2) by building a theoretical guarantee on ScD’s consistency for any distribution $\mathcal{F}$ satisfying Equation (6) under the DCMMDF model. Set $\tau=\mathrm{max}_{i,j\in[n]}|A(i,j)-\Omega(i,j)|$ and $\gamma=\mathrm{max}_{i,j\in[n]}\frac{\mathrm{Var}(A(i,j))}{\theta(i)\theta(j)}$, where $\mathrm{Var}(A(i,j))$ denotes the variance of $A(i,j)$. $\tau$ and $\gamma$ are two parameters related to distribution $\mathcal{F}$. For different distributions $\mathcal{F}$, the upper bounds of $\tau$ and $\gamma$ can be different, see Examples 1-5 for detail. We need the following assumption for our theoretical analysis.
@@
Under $DCMMDF_{n}(K,P,\Pi,\Theta,\mathcal{F})$, assume $\gamma\theta_{\mathrm{max}}\|\theta\|_{1}\geq\tau^{2}\mathrm{log}(n)$.
Assumption 1 provides a lower bound requirement on $\gamma\theta_{\mathrm{max}}\|\theta\|_{1}$ for our theoretical analysis and it functions similar to Assumption 3.1 of mao2020estimating and Assumption 1 of MixedSCORE . We also need the following condition which simplifies our theoretical result.
@@
Under $DCMMDF_{n}(K,P,\Pi,\Theta,\mathcal{F})$, assume $\kappa(P)=O(1),K=O(1),\lambda_{K}(\Pi^{\prime}\Pi)=O(\frac{n}{K}),\theta_{\mathrm{max}}=O(\sqrt{\rho})$, and $\theta_{\mathrm{min}}=O(\sqrt{\rho})$ for some $\rho>0$.
In Condition 1, $\kappa(P)=O(1)$ is mild because it means the connectivity matrix $P$ is well-conditioned, $K=O(1)$ means we only consider a constant number of communities for our theoretical analysis, $\lambda_{K}(\Pi^{\prime}\Pi)=O(\frac{n}{K})$ means the size of each community is in the same order, and $\theta_{\mathrm{max}}=O(\sqrt{\rho}),\theta_{\mathrm{min}}=O(\sqrt{\rho})$ mean that all entries of $\theta$ are in the same order. So, we see that Condition 1 is mild. We are ready to present the main theorem:
@@
Under $DCMMDF_{n}(K,P,\Pi,\Theta,\mathcal{F})$, let $\hat{\Pi}$ be obtained from Algorithm 1, when Assumption 1 and Condition 1 hold, there exists a permutation matrix $\mathcal{P}\in\mathbb{R}^{K\times K}$ such that with probability at least $1-o(n^{-5})$, we have
@@
Proof of Theorem 1 is provided in A. In Theorem 1, when $P,K,\theta_{\mathrm{min}}$, and distribution $\mathcal{F}$ are fixed, increasing $n$ decreases ScD’s error rate. In particular, when $n$ goes to infinity, the theoretical upper bound of the error rate in Theorem 1 goes to zero, and this guarantees ScD’s consistency under our DCMMDF. In Theorem 1, only parameters $\gamma$ and $\tau$ are directly related to distribution $\mathcal{F}$, so our main theoretical result in Theorem 1 is general since we do not constrain $A$’s distribution $\mathcal{F}$. In the following examples, we will show that $\gamma$ is finite and we can obtain a more accurate theoretical upper bound of ScD’s error rate by analyzing $\gamma$’s upper bound for some specific distributions.
@@
When $\mathcal{F}$ is Normal distribution such that $A(i,j)\sim\mathrm{Normal}(\Omega(i,j),\sigma^{2}_{A})$, i.e., $A(i,j)\in\mathbb{R}$, we have $\mathbb{E}[A(i,j)]=\Omega(i,j)$ satisfying Equation (6), $P$’s non-diagonal elements can be any finite real values because the mean of Normal distribution can be negative, $\Omega(i,j)$’s range is (,)($\infty$,$\infty$)( - ∞ , ∞ ), $\tau$ is an unknown finite positive real value, and $\gamma=\mathrm{max}_{i,j\in[n]}\frac{\mathrm{Var}(A(i,j))}{\theta(i)\theta(j)}\leq\frac{\sigma^{2}_{A}}{\theta^{2}_{\mathrm{min}}}=O(\frac{\sigma^{2}_{A}}{\rho})$ is finite. Setting $\gamma=O(\frac{\sigma^{2}_{A}}{\rho})$ in Theorem 1 obtains the theoretical upper bound of ScD’s error rate when $\mathcal{F}$ is Normal distribution. We see that increasing $\rho$ (or decreasing $\sigma^{2}_{A}$) decreases ScD’s error rate for Normal distribution.
@@
When $\mathcal{F}$ is Bernoulli distribution such that $A(i,j)\sim\mathrm{Bernoulli}(\Omega(i,j))$ and our DCMMDF degenerates to the DCMM model, i.e., $A(i,j)\in\{0,1\}$, we have $\mathbb{E}[A(i,j)]=\Omega(i,j)$ satisfying Equation (6), all entries of $P$ should be nonnegative, $\Omega(i,j)$ is a probability in $[0,1]$, and $\tau=1$. For this case, we have $\mathrm{Var}(A(i,j))=\Omega(i,j)(1-\Omega(i,j))\leq\Omega(i,j)=\theta(i)\theta(j)\Pi(i,:)P\Pi^{\prime}(j,:)\leq\theta(i)\theta(j)P_{\mathrm{max}}$, which suggests that $\gamma=P_{\mathrm{max}}$, a finite number. Setting $\gamma=P_{\mathrm{max}}$ in Theorem 1 gets the theoretical upper bound of ScD’s error rate and we find that increasing $\rho$ decreases ScD’s error rate. Setting $\gamma=P_{\mathrm{max}}$ and $\tau=1$, Assumption 1 is $\rho n\geq O(\mathrm{log}(n))$ by Condition 1, and it means a lower bound requirement on network sparsity. So $\rho$ controls network sparsity for Bernoulli distribution.
@@
When $\mathcal{F}$ is Poisson distribution such that $A(i,j)\sim\mathrm{Poisson}(\Omega(i,j))$, i.e., $A(i,j)\in\mathbb{N}$, where $\mathbb{N}$ is the set of nonnegative integers, we have $\mathbb{E}[A(i,j)]=\Omega(i,j)$ satisfying Equation (6), $P$’s elements should be nonnegative, $\Omega(i,j)$’s range is $[0,\infty)$, $\tau$ is an unknown positive integer, and $\gamma=\mathrm{max}_{i,j\in[n]}\frac{\mathrm{Var}(A(i,j))}{\theta(i)\theta(j)}=\mathrm{max}_{i,j\in[n]}\frac{\Omega(i,j)}{\theta(i)\theta(j)}\leq P_{\mathrm{max}}$, i.e., $\gamma$ is finite. Setting $\gamma=P_{\mathrm{max}}$ in Theorem 1 gets the theoretical result for Poisson distribution and we find that increasing $\rho$ decreases ScD’s error rate.
@@
When $\mathcal{F}$ is Uniform distribution such that $A(i,j)\sim\mathrm{Uniform}(0,2\Omega(i,j))$ for $i,j\in[n]$, i.e., $A(i,j)\in(0,2\rho)$, we have $\mathbb{E}[A(i,j)]=\frac{0+2\Omega(i,j)}{2}=\Omega(i,j)$ satisfying Equation (6), $P$’s entries should be nonnegative, $\Omega(i,j)$’s range is $[0,+\infty)$, $\tau$ is an unknown value with a upper bound $2\rho$, and $\gamma=\mathrm{max}_{i,j\in[n]}\frac{\mathbb{E}[(A(i,j)-\Omega(i,j))^{2}]}{\theta(i)\theta(j)}=\mathrm{max}_{i,j\in[n]}\frac{4\Omega^{2}(i,j)}{12\theta(i)\theta(j)}\leq\mathrm{max}_{i,j\in[n]}\frac{(\theta(i)\theta(j)P_{\mathrm{max}})^{2}}{3\theta(i)\theta(j)}=O(\frac{\rho P^{2}_{\mathrm{max}}}{3})$. Setting $\gamma=O(\frac{\rho P^{2}_{\mathrm{max}}}{3})$ in Theorem 1, we find that $\rho$ disappears in the theoretical upper bound of ScD’s error rate, which indicates that increasing $\rho$ has almost no significant impact on the behavior of ScD.
@@
DCMMDF can also model overlapping signed networks by setting $\mathbb{P}(A(i,j)=1)=\frac{1+\Omega(i,j)}{2}$ and $\mathbb{P}(A(i,j)=-1)=\frac{1-\Omega(i,j)}{2}$, i.e., $A(i,j)\in\{1,-1\}$ for $i\neq j$. For signed networks, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds by basic computation, $P$’s non-diagonal elements can be negative, $\Omega(i,j)$’s range is $[-1,1]$ because $\frac{1+\Omega(i,j)}{2}$ and $\frac{1-\Omega(i,j)}{2}$ are probabilities, $\tau$ is $2$, and $\mathrm{Var}(A(i,j))=1-\Omega^{2}(i,j)\leq 1$, i.e., $\gamma\leq\frac{1}{\theta^{2}_{\mathrm{min}}}=O(\frac{1}{\rho})$ is finite. Setting $\gamma=O(\frac{1}{\rho})$ in Theorem 1 obtains theoretical results for the signed network and we see that increasing $\rho$ decreases ScD’s error rate.
Generally speaking, benefiting from the distribution-free property of our DCMMDF, $\mathcal{F}$ can be any distribution as long as Equation (6) holds such as the Binomial, Double exponential, Gamma, Exponential, and Laplace distributions in http://www.stat.rice.edu/~dobelman/courses/texts/distributions.c&b.pdf.
@@
By Examples 1, 4, and 5, we find that the non-diagonal elements of $A$ are always nonzero, which suggests that there is always a link between any two nodes. However, this is impractical for real-world large-scale networks in which most node pairs have no connections lei2015consistency . Treat an edge with weight 0 as a missing edge. Following the idea in qing2021DFM , we can generate missing edges for networks generated from DCMMDF such that $A(i,j)$ is multiplied by $\mathcal{A}(i,j)$ for $i,j\in[n]$, where $\mathcal{A}\in\{0,1\}^{n\times n}$ is a symmetric and connected adjacency matrix for an un-weighted network and it can be generated from any model for un-directed and un-weighted networks such as SBM, DCSBM, MMSB, DCMM, and Erdös-Rényi random graph.

### v3

@@ 4 Asymptotic Consistency
In this section, we aim to answer Q (2) by building a theoretical guarantee on ScD’s consistency for any distribution $\mathcal{F}$ satisfying Equation (6) under the WDCMM model. Set $\tau=\mathrm{max}_{i,j\in[n]}|A(i,j)-\Omega(i,j)|$ and $\gamma=\mathrm{max}_{i,j\in[n]}\frac{\mathrm{Var}(A(i,j))}{\theta(i)\theta(j)}$, where $\mathrm{Var}(A(i,j))=\mathbb{E}[(A(i,j)-\Omega(i,j))^{2}]$ denotes the variance of $A(i,j)$. $\tau$ and $\gamma$ are two parameters related to distribution $\mathcal{F}$. For different distributions $\mathcal{F}$, the upper bounds of $\tau$ and $\gamma$ can be different, see Examples 1-5 for detail. We need the following assumption for our theoretical analysis.
@@
Under $WDCMM_{n}(K,P,\Pi,\Theta,\mathcal{F})$, assume that $\tau$ is finite or at least finite with high probability.
In this paper, we do not require $\tau$ to be a constant; rather, we assume that $\tau$ is bounded above by some quantity (which may depend on $n$) with high probability. This is sufficient for applying the row-wise eigenspace error bound from Theorem 4.2.1 of [66], which only requires that the entries of $A-\Omega$ are uniformly bounded in absolute value by some $\tau$ (not necessarily constant). Although $\tau$ may grow with $n$ for certain distributions (e.g., Normal and Poisson), Assumption 1 ensures that it is finite with high probability. Indeed, Lemmas 1 and 2 in Appendix B establish that $\tau$ is bounded with high probability for Normal and Poisson distributions, respectively, while Lemma 3 provides a general finite-with-high-probability bound of $\tau$ for any distribution with finite variance.
Assumption 2.
Under $WDCMM_{n}(K,P,\Pi,\Theta,\mathcal{F})$, assume $\gamma\theta_{\mathrm{max}}\|\theta\|_{1}\geq\tau^{2}\mathrm{log}(n)$.
Assumption 2 provides a lower bound requirement on $\gamma\theta_{\mathrm{max}}\|\theta\|_{1}$ for our theoretical analysis and it functions similar to Assumption 3.1 of [36] and Assumption 1 of [33]. Note that Assumption 2 presumes the fulfillment of Assumption 1, implying that $\tau$ can be regarded as a constant. We also need the following condition which simplifies our theoretical result.
@@
Under $WDCMM_{n}(K,P,\Pi,\Theta,\mathcal{F})$, assume $\kappa(P)=O(1),K=O(1),\lambda_{K}(\Pi^{\prime}\Pi)=O(\frac{n}{K}),\theta_{\mathrm{max}}=O(\sqrt{\rho})$, and $\theta_{\mathrm{min}}=O(\sqrt{\rho})$ for some $\rho>0$.
In Condition 1, $\kappa(P)=O(1)$ is mild because it means the connectivity matrix $P$ is well-conditioned, $K=O(1)$ means we only consider a constant number of communities for our theoretical analysis, $\lambda_{K}(\Pi^{\prime}\Pi)=O(\frac{n}{K})$ means the size of each community is in the same order, and $\theta_{\mathrm{max}}=O(\sqrt{\rho}),\theta_{\mathrm{min}}=O(\sqrt{\rho})$ mean that all entries of $\theta$ are in the same order. So, we see that Condition 1 is mild. Recall that when $\mathcal{F}$ is Bernoulli distribution ( and all nodes are pure), and $\theta(i)=\sqrt{\rho}$ for $i\in[n]$, WDCMM reduces to the classical MMSB (and SBM) model. For this case, the sparsity parameter $\rho$ governs network’s sparsity [24, 36]. Meanwhile, Corollary 3.1 in [36] and Assumptions 1-3 in [33] also adopt similar conditions to streamline their theoretical analysis. We are ready to present the main theorem:
@@
Under $WDCMM_{n}(K,P,\Pi,\Theta,\mathcal{F})$, let $\hat{\Pi}$ be obtained from Algorithm 1, when Assumption 1, Assumption 2, and Condition 1 hold, there exists a permutation matrix $\mathcal{P}\in\mathbb{R}^{K\times K}$ such that with probability at least $1-o(n^{-5})$, we have
@@
Proof of Theorem 1 is provided in Appendix A. In Theorem 1, when $P,K,\theta_{\mathrm{min}}$, and distribution $\mathcal{F}$ are fixed, increasing $n$ decreases ScD’s error rate. In particular, when $n$ goes to infinity, the theoretical upper bound of the error rate in Theorem 1 goes to zero, and this guarantees ScD’s consistency under our WDCMM. In Theorem 1, only parameters $\gamma$ and $\tau$ are directly related to distribution $\mathcal{F}$, so our main theoretical result in Theorem 1 is general since we do not constrain $A$’s distribution $\mathcal{F}$. In the following examples, we will show that Assumption 1 is satisfied, $\gamma$ is finite, and we can obtain a more accurate theoretical upper bound of ScD’s error rate by analyzing $\gamma$’s upper bound for some specific distributions.
@@
When $\mathcal{F}$ is Normal distribution such that $A(i,j)\sim\mathrm{Normal}(\Omega(i,j),\sigma^{2}_{A})$ for $\sigma_{A}>0$, i.e., $A(i,j)\in\mathbb{R}$, we have $\mathbb{E}[A(i,j)]=\Omega(i,j)$ satisfying Equation (6), $P$’s non-diagonal elements can be any finite real values because the mean of Normal distribution can be negative, $\Omega(i,j)$’s range is $(-\infty,\infty)$, $\tau$ is finite with high probability by Lemma 1 in Appendix B, and $\gamma=\mathrm{max}_{i,j\in[n]}\frac{\mathrm{Var}(A(i,j))}{\theta(i)\theta(j)}\leq\frac{\sigma^{2}_{A}}{\theta^{2}_{\mathrm{min}}}=O(\frac{\sigma^{2}_{A}}{\rho})$ is finite. Setting $\gamma=O(\frac{\sigma^{2}_{A}}{\rho})$ in Theorem 1 obtains $O(\frac{\sigma_{A}\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\rho\sqrt{n}})$, the theoretical upper bound of ScD’s error rate when $\mathcal{F}$ is Normal distribution. We see that increasing $\rho$ (or decreasing $\sigma^{2}_{A}$) decreases ScD’s error rate for Normal distribution.
@@
When $\mathcal{F}$ is Bernoulli distribution such that $A(i,j)\sim\mathrm{Bernoulli}(\Omega(i,j))$ and our WDCMM degenerates to the DCMM model, i.e., $A(i,j)\in\{0,1\}$, we have $\mathbb{E}[A(i,j)]=\Omega(i,j)$ satisfying Equation (6), all entries of $P$ should be nonnegative, $\Omega(i,j)$ is a probability in $[0,1]$, and $\tau=1$. For this case, we have $\mathrm{Var}(A(i,j))=\Omega(i,j)(1-\Omega(i,j))\leq\Omega(i,j)=\theta(i)\theta(j)\Pi(i,:)P\Pi^{\prime}(j,:)\leq\theta(i)\theta(j)P_{\mathrm{max}}$, which suggests that $\gamma=P_{\mathrm{max}}$, a finite number. Setting $\gamma=P_{\mathrm{max}}$ in Theorem 1 gets $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{\rho n}})$, the theoretical upper bound of ScD’s error rate. We find that increasing $\rho$ decreases ScD’s error rate. Setting $\gamma=P_{\mathrm{max}}$ and $\tau=1$, Assumption 2 is $\rho n\geq O(\mathrm{log}(n))$ by Condition 1, and it means a lower bound requirement on network sparsity. So $\rho$ controls network sparsity for Bernoulli distribution.
Remark 7.
When $\mathcal{F}$ is Bernoulli distribution, we observe that our sparsity requirement, $\rho n\geq O(\mathrm{log}(n))$, aligns with the sparsity conditions stated in Theorem 3.1 of [24], Equation (2.9) of [25], Equation (2.8) in Assumption 2 of [67], and Assumption 1 of [33]. Notably, such sparsity condition is mild as it merely necessitate that the average node degree increases at a rate faster than $\mathrm{log}(n)$. Meanwhile, our error bound $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{\rho n}})$ also matches with that in Equation (F.76) of [33] or the one in Theorem 2.2 in the latest arXiv version of [33] when Condition 1 holds. This indicates that our proposed method achieves a comparable theoretical guarantee in terms of error rate. Additionally, by the analysis in Remarks 5, 11, and 13 of [62], we know that the error bounds in Theorem 3.2 of [36] and Theorem 3.2 of [35] should multiply a $\log^{\xi}(n)$ term for some constant $\xi>1$. After multiplying this term, their error bounds become $O\left(\frac{\log^{\xi}(n)}{\sigma_{K}(P)\sqrt{\rho n}}\right)$, which is larger than ours.
@@
When $\mathcal{F}$ is Poisson distribution such that $A(i,j)\sim\mathrm{Poisson}(\Omega(i,j))$, i.e., $A(i,j)\in\mathbb{N}$, we have $\mathbb{E}[A(i,j)]=\Omega(i,j)$ satisfying Equation (6), $P$’s elements should be nonnegative, $\Omega(i,j)$’s range is $0,\infty)$, $\tau$ is finite with high probability by Lemma 2 in Appendix B, and $\gamma=\mathrm{max}_{i,j\in[n]}\frac{\mathrm{Var}(A(i,j))}{\theta(i)\theta(j)}=\mathrm{max}_{i,j\in[n]}\frac{\Omega(i,j)}{\theta(i)\theta(j)}\leq P_{\mathrm{max}}$, i.e., $\gamma$ is finite. Setting $\gamma=P_{\mathrm{max}}$ in Theorem 1 gets $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{\rho n}})$, the theoretical result for Poisson distribution. We find that increasing $\rho$ decreases ScD’s error rate.
@@
When $\mathcal{F}$ is Uniform distribution such that $A(i,j)\sim\mathrm{Uniform}(0,2\Omega(i,j))$ for $i,j\in[n]$, i.e., $A(i,j)\in(0,2\Omega(i,j))$, we have $\mathbb{E}[A(i,j)]=\frac{0+2\Omega(i,j)}{2}=\Omega(i,j)$ satisfying Equation (6), $P$’s entries should be nonnegative, $\Omega(i,j)$’s range is $0,+\infty)$, $\tau$’s upper bound is $2\theta^{2}_{\mathrm{max}}P_{\mathrm{max}}$, and $\gamma=\mathrm{max}_{i,j\in[n]}\frac{\mathbb{E}[(A(i,j)-\Omega(i,j))^{2}]}{\theta(i)\theta(j)}=\mathrm{max}_{i,j\in[n]}\frac{4\Omega^{2}(i,j)}{12\theta(i)\theta(j)}\leq\mathrm{max}_{i,j\in[n]}\frac{(\theta(i)\theta(j)P_{\mathrm{max}})^{2}}{3\theta(i)\theta(j)}=O(\frac{\rho P^{2}_{\mathrm{max}}}{3})$. Setting $\gamma=O(\frac{\rho P^{2}_{\mathrm{max}}}{3})$ in Theorem 1, we find that $\rho$ disappears in $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{n}})$ (the theoretical upper bound of ScD’s error rate), which indicates that increasing $\rho$ has almost no significant impact on the behavior of ScD.
@@
WDCMM can also model overlapping signed networks by setting $\mathbb{P}(A(i,j)=1)=\frac{1+\Omega(i,j)}{2}$ and $\mathbb{P}(A(i,j)=-1)=\frac{1-\Omega(i,j)}{2}$, i.e., $A(i,j)\in\{1,-1\}$ for $i\neq j$. For signed networks, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds by basic computation, $P$’s non-diagonal elements can be negative, $\Omega(i,j)$’s range is [1,1][-1,1] because $\frac{1+\Omega(i,j)}{2}$ and $\frac{1-\Omega(i,j)}{2}$ are probabilities, $\tau$ is $2$, and $\mathrm{Var}(A(i,j))=1-\Omega^{2}(i,j)\leq 1$, i.e., $\gamma\leq\frac{1}{\theta^{2}_{\mathrm{min}}}=O(\frac{1}{\rho})$ is finite. Setting $\gamma=O(\frac{1}{\rho})$ in Theorem 1 obtains $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\rho\sqrt{n}})$, the theoretical results for signed networks. We see that increasing $\rho$ decreases ScD’s error rate.
Generally speaking, benefiting from the distribution-free property of our WDCMM, $\mathcal{F}$ can be any distribution as long as Equation (6) holds, as previously discussed in the analysis following Definition 1. Furthermore, by adopting a similar approach (i.e., analyzing $\tau$’s finiteness and $\gamma$’s upper bound under different distributions) to the analyses presented in Examples 1-5, we can delve deeper into the theoretical upper bound of ScD’s error rate for various other distributions listed after Definition 1. For brevity, we list these theoretical results for different distributions in Table 2. Following a similar analysis to Examples 1-5, we can analyze ScD’s performance under different distributions. To save space, we omit the detailed analysis here.
Distribution $\mathcal{F}$ $A(i,j)$’s range $\Omega(i,j)$’s range $\tau$’s upper bound $\gamma$’s upper bound Theoretical upper bound of ScD’ error rate
$\mathrm{Normal}(\Omega(i,j),\sigma^{2}_{A})$ for $\sigma_{A}>0$ $\mathbb{R}$ $\mathbb{R}$ FWHP $O(\frac{\sigma^{2}_{A}}{\rho})$ $O(\frac{\sigma_{A}\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\rho\sqrt{n}})$
$\mathrm{Bernoulli}(\Omega(i,j))$ $\{0,1\}$ $[0,1]$ 1 $P_{\mathrm{max}}$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{\rho n}})$
$\mathrm{Poisson}(\Omega(i,j))$ $\mathbb{N}$ $\mathbb{R}_{+}$ FWHP $P_{\mathrm{max}}$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{\rho n}})$
$\mathrm{Uniform}(0,2\Omega(i,j))$ $(0,2\Omega(i,j))$ $\mathbb{R}_{+}$ $2\theta^{2}_{\mathrm{max}}P_{\mathrm{max}}$ $O(\frac{\rho P^{2}_{\mathrm{max}}}{3})$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{n}})$
Signed network $\{-1,1\}$ [1,1][-1,1] 2 $O(\frac{1}{\rho})$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\rho\sqrt{n}})$
$\mathrm{Binomial}(m,\frac{\Omega(i,j)}{m})$ $\{0,1,2\ldots,m\}$ $[0,m]$ $m$ $P_{\mathrm{max}}$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{\rho n}})$
$\mathrm{Negative~binomial}(m,\frac{m}{m+\Omega(i,j)})$ $\mathbb{N}$ $\mathbb{R}_{+}$ FWHP $O(\frac{\rho P^{2}_{\mathrm{max}}}{m})$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{n}})$
$\mathrm{Exponential}(\frac{1}{\Omega(i,j)})$ $\mathbb{R}_{+}$ $\mathbb{R}_{>0}$ FWHP $O(\rho P^{2}_{\mathrm{max}})$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{n}})$
$\mathrm{Gamma}(a,\frac{\Omega(i,j)}{a})$ for $a>0$ $\mathbb{R}_{>0}$ $\mathbb{R}_{>0}$ FWHP $O(\frac{\rho P^{2}_{\mathrm{max}}}{a})$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{n}})$
$\mathrm{Gamma}(\frac{\Omega(i,j)}{a},a)$ for $a>0$ $\mathbb{R}_{>0}$ $\mathbb{R}_{>0}$ FWHP $aP_{\mathrm{max}}$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{\rho n}})$
$\mathrm{Beta}(\frac{a\Omega(i,j)}{1-\Omega(i,j)},a)$ for $a>0$ $(0,1)$ $(0,1)$ 1 $\frac{P_{\mathrm{max}}}{a}$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{\rho n}})$
$\mathrm{Beta}(a,\frac{a(1-\Omega(i,j))}{\Omega(i,j)})$ for $a>0$ $(0,1)$ $(0,1)$ 1 $O(\frac{\rho P^{2}_{\mathrm{max}}}{a})$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{n}})$
$\mathrm{Geometric}(\frac{1}{\Omega(i,j)})$ $\mathbb{N}_{+}$ $1,+\infty)$ FWHP $O(\rho P^{2}_{\mathrm{max}})$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{n}})$
$\mathrm{Laplace}(\Omega(i,j),a)$ for $a>0$ $\mathbb{R}$ $\mathbb{R}$ FWHP $O(\frac{a^{2}}{\rho})$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\rho\sqrt{n}})$
$\mathrm{Lognormal}(\mathrm{log}(\Omega(i,j))-\frac{a^{2}}{2},a^{2})$ for $a>0$ $\mathbb{R}_{>0}$ $\mathbb{R}_{>0}$ FWHP $O(\rho P^{2}_{\mathrm{max}}(\mathrm{exp}(a^{2})-1))$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{n}})$
$\mathrm{Pareto}(\frac{a-1}{a}\Omega(i,j),a)$ for $a>2$ $(\frac{a-1}{a}\Omega(i,j),+\infty)$ $\mathbb{R}_{>0}$ FWHP $O(\frac{\rho P^{2}_{\mathrm{max}}}{a(a-2)})$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{n}})$
$\mathrm{Logistic}(\Omega(i,j),a)$ for $a>0$ $\mathbb{R}$ $\mathbb{R}$ FWHP $O(\frac{a^{2}}{\rho})$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\rho\sqrt{n}})$
$\mathrm{Weibull}(a,(\frac{\Omega(i,j)}{\Gamma(1+\frac{1}{a})})^{a})$ for $a>0$ $\mathbb{R}_{+}$ $\mathbb{R}_{>0}$ FWHP $O(\frac{\rho P^{2}_{\mathrm{max}}\Gamma(1+\frac{2}{a})}{\Gamma^{2}(1+\frac{1}{a})})$ $O(\frac{\sqrt{\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{n}})$
Table 2: Theoretical results for different distributions under the WDCMM model, where FWHP stands for “finite with high probability”. Here, $i$ and $j$ range in $[n]$. The FWHP property of $\tau$ is guaranteed by Lemma 1 for the Normal distribution, Lemma 2 for the Poisson distribution, and Lemma 3 for other distributions, as detailed in Appendix B.
@@
By Examples 1, 4, and 5, we find that the non-diagonal elements of $A$ are always nonzero, which suggests that there is always a link between any two nodes. However, this is impractical for real-world large-scale networks in which most node pairs have no connections [24]. Treat an edge with weight 0 as a missing edge. We can generate missing edges for networks generated from WDCMM such that $A(i,j)$ is multiplied by $\mathcal{A}(i,j)$ for $i,j\in[n]$, where $\mathcal{A}\in\{0,1\}^{n\times n}$ is a symmetric and connected adjacency matrix for an un-weighted network and it can be generated from any model for un-directed and un-weighted networks such as SBM, DCSBM, MMSB, DCMM, and Erdös-Rényi random graph.